# Two-stage Generative Question Answering on Temporal Knowledge Graph Using Large Language Models

### v1

@@ 5.1 Datasets, Metrics and Baselines
CronQuestions Saxena et al. (2021) is a temporal QA dataset based on the Wikidata TKG Lacroix et al. (2020), comprising 125k entities, 203 relations, 1.7k timestamps and 328k facts. It contains 410K unique question-answer pairs, including annotated entities and timestamps, with 350k for training and 30k for validation and testing. The dataset can be categorized into simple reasoning (Simple Entity and Simple Time) and complex reasoning (Before/After, First/Last, and Time Join) based on temporal constraints. Following previous studies, we use two popular evaluation metrics, Hits@1 and Hits@10. More information about datasets and metrics can be found in Appendix A.
We compare four types of baselines: 1) KG embedding-based models including EaE Feng et al. (2020b) and EmbedKGQA Saxena et al. (2020); 2) TKG embedding-based models including CronKGQA Saxena et al. (2021), EntityQR Mavromatis et al. (2022), TMA Liu et al. (2023a), TSQA Shang et al. (2022) and TempoQA Mavromatis et al. (2022); and 3) Language models, including BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and ChatGPT OpenAI (2023). The implementation details of the baselines and our model are described in Appendix B.
@@ 5.2 Main Results
Table 1 reports the performance of all methods on the CronQuestions dataset for various question types. We can observe that GenTKGQA consistently outperforms the baselines in terms of "Overall" performance, and achieves significant improvements of 11.3% in the "Complex" question type and 10.6% in the "Time" answer type on the Hits@1 metric over the second best method. Especially, our model achieves 100% for the "Simple" question type. The possible reason is that "Simple" questions usually involve single facts, GenTKGQA can easily retrieve the relevant facts containing the answer and infer the correct answer through instruction tuning technique.
Furthermore, compared to KG embedding methods, temporal KG embedding methods show significant results on various metrics, thanks to the fact that the temporal information of the TKG is taken into account in the question representation. This is also why KG embedding methods are particularly ineffective for the "Time" answer type. However, most TKG embedding methods treat the QA task as a link prediction, which works for the "Simple" question type containing a single fact compared to the "Complex".
We find that PLMs (BERT, RoBERTa) and LLMs (ChatGPT) have the lowest performance on the TKGQA task without TKG information. This suggests that language models (LM), whether encoded or generated, with a large or small number of parameters, have difficulty answering temporal questions without any relevant context. w/ tkg indicates that LMs use entity/time embeddings or relevant facts from the TKG. Obviously, LMs w/ tkg have significantly better performance, which suggests that the language models have some degree of temporal reasoning capability when relevant TKG information is provided, validating the importance of the subgraph retrieval phase. It is worth noting that ChatGPT w/ tkg still performs weakly in reasoning about complex problem types when providing the facts retrieved in the first stage by GenTKGQA, while our model achieves the best results. This demonstrates the effectiveness of interacting GNN and LM representations in a non-shallow way in dealing with complex temporal problems. The above findings demonstrate the adequacy of our two motivations for solving complex temporal problems with LLMs. A possible reason for the poor improvement of our model’s Hits@10 metric for the "Entity" type is that the generative LMs generate irrelevant responses when asked to generate multiple answers. Last but not least, GenTKGQA, as a generative QA model, achieves better results than traditional extractive QA methods.
@@
5.3 Ablation Study
w/o SR means that we directly perform problem inference without using relevant subgraph information in both the model training and inference phases, while w/o SR inference means that we do not provide subgraphs only at the inference. We can observe a sharp decrease in model effectiveness due to the lack of use of subgraph information, which is consistent with the other LMs (Section 5.2). This result shows that the current LMs do not have the reasoning capability to deal directly with temporal reasoning problems, validating the importance of the subgraph retrieval module. The w/o SR inference result indicates that GenTKGQA remembers part of the structured knowledge during the training phase and improves the temporal inference performance without providing subgraph information. w/ SR random denotes the random selection of relevant facts involving entities in the question, and the drop in results proves the validity of our first-stage approach.
w/o T-GNN indicates that we directly use temporal embeddings to represent entities and relations of subgraphs. The slight decrease in the results indicates that the T-GNN is able to perceive the structural information of the TKG. w/o VKI means that we try to remove the virtual knowledge indicators from the input prompt. Model performance degradation shows that indicators can bridge the gap between distinct representations.
Figure 3: Parameter sensitivity on our GenTKGQA.
@@ 5.3 Ablation Study
Before/ What was the first General Inspector
After of the Bundeswehr after 2004?
[Wolfgang Schneiderhan, position held, Bundeswehr Chief, 2002, 2009] [Volker Wieker] [Wolfgang Schneiderhan]
[Volker Wieker, position held, Bundeswehr Chief, 2010, 2018]
@@
Time Who worked with Christiane Hoffmann
Join during his employment with FAZ? [Gero von Randow, Frank Schirrmacher,
[Christiane Hoffmann, employer, FAZ, 1994, 2010] [Christiane Hoffmann] Oliver Gehrs, Peter Kiefer,
[Gero von Randow, employer, FAZ, 2001, 2003] Wolfgang Bresser, Ulrich Kuehl
[Frank Schirrmacher, employer, FAZ, 1994, 2014] Henning Franke, Stephan Löwenstein]
@@ 5.4 Sensitivity Analysis
We explore the impact of different training data sizes to reason about complex temporal questions. As shown in Figure 3(a), by comparing the Hits@1 metric of several methods for the complex question type, our method consistently outperforms others as the training data expands. In particular, at 20% of the training data, our model outperforms the second best model by 32.3%, demonstrating that our model has strong inference ability in the case of few-shot samples due to its intrinsic knowledge.
@@
We report the performance changes on the CronQuestions dataset by varying the number of retrieved facts $n$ in Figure 3(b). It can be seen that the model performs poorly with a small number of relevant facts ($n$=3), and there is a slight drop in performance at $n$=8. Fewer facts do not provide sufficient context knowledge, while more facts may introduce noise. Taking this into consideration, we set the hyper-parameter $n$ to 5.
@@ 5.5 Qualitative Results
When providing relevant facts retrieved by GenTKGQA as background knowledge, ChatGPT performs competitively in the simple question type, correctly answering questions with entity or time as the answer. However, it has difficulty answering complex types of questions. For example, in the "Before/After" and "First/Last" question types, ChatGPT struggles to understand the temporal order of the relevant facts in the time dimension and gives incorrect answers. Besides, it does not seem to understand the dependencies among related facts. In the case of the "Time Join" question type, which asks for other entities who worked with Christiane Hoffmann at a certain time, ChatGPT wrongly considers the entity Christiane Hoffmann itself that appears in the relevant facts as the correct answer. On the contrary, GenTKGQA performs well in both simple and complex question types, thanks to the fact that our proposed approach allows the LLM to perceive the structural and temporal information of the retrieved subgraphs. However, similar to other generative LLMs, GenTKGQA randomly generates some irrelevant answers when generating multiple answers, e.g., in the "Simple Entity" and "Time Join" question types.
@@ 6 Conclusion
We propose a novel generative framework, GenTKGQA, which guides the LLM in a two-stage manner to handle temporal question answering on TKG. Specifically, at the subgraph retrieval phase, we adopt a divide-and-conquer strategy to exploit the LLM’s intrinsic knowledge to mine the temporal constraints and structural links in the temporal questions to reduce the search space of the subgraphs in both time-space dimensions. We employ the in-context learning approach to complete subgraph retrieval for the entire dataset with a small number of samples. In order to improve the inference performance of the LLM on complex types of problems, at the answer generation phase, we present the instruction tuning technique to make the open-source LLM truly understand the temporal order and structural dependencies among retrieved facts. Most critically, we design novel knowledge indicators to establish a bridge between subgraph neural information and text representations. Experimental results show that our framework can effectively utilize the LLM to solve the complex question type of TKGQA task and validate the adequacy of our motivation.
@@
Although the complex temporal question in the CronQuestions dataset involves multiple facts, the inter-entity connection in each fact is single-hop, so the hyper-parameter $k$ of our model is set to 1 to achieve the best results. In fact, the vast majority of current TKGQA datasets involve facts that are single-hop. So, in the future, we will explore more datasets to solve inference for multi-hop complex temporal problems over TKG.
@@ Appendix A Dataset Statistics and Metrics
We use the CronQuestions dataset in our experiments. Dataset statistics and TKG information are described in Table 4 and 5, respectively.
@@ Appendix B Baselines and Implementation Details
We use the OpenAI-API [^1] https://platform.openai.com/docs/api-reference (gpt-3.5-turbo-0613222https://platform.openai.com/docs/models/gpt-3-5-turbo) for all ChatGPT-related experiments, including subsequent ChatGPT baselines.
In the subgraph retrieval phase, we use ChatGPT OpenAI (2023) to mine temporal constraints and structural links between entities and add 5 samples to the in-context learning prompt templates, which are presented as Table 6 and 7. We set $k$=1 for the top-$k$ relations. In the answer generation phase, following Lacroix et al. (2020), we select the dimension of entity/relation/time embeddings to 512. For T-GNN, the layer $l$ is set to 1, the linear transformations $\mathbf{W}_{q}$, $\mathbf{W}_{k}$ and $\mathbf{W}_{m}$ are 512×512, and the $\mathbf{m}$ and $\mathbf{b}$ of the MLP layer are $|\mathcal{E}|$. We use the open-source Llama 2-7B Touvron et al. (2023) for instruction tuning and select up to $n$=5 relevant facts as additional knowledge. The linear projection layer $\mathbf{W}_{p}$ is 512×4096. We fine-tune Llama 2-7B using LoRA Hu et al. (2022) with rank 64. The number of epochs is set to 3 and the learning rate is 3e-4. We use the AdamW optimizer Loshchilov and Hutter (2019) with a fixed batch size of 8. We conduct all the experiments with NVIDIA A100 GPUs, and the results of each experiment are averaged over three runs. We will release the source code upon acceptance.
@@
TSQR Chen et al. (2021b): TSQA presents a contrastive learning module that improves sensitivity to time relation words.
@@
ChatGPT OpenAI (2023): To ensure that the output format meets the expected requirements, we use the in-context learning approach to motivate ChatGPT to answer the questions and provide 5 examples in the prompt template. The specific templates are presented in Table 9 of Appendix C. w/o tkg and w/ tkg differ in whether or not question-relevant facts are provided in the input prompts, which are retrieved in the first stage by our proposed GenTKGQA framework.

### v2

@@ 5.1 Datasets, Metrics and Baselines
CronQuestions Saxena et al. (2021) is a temporal QA dataset, which contains 410K unique question-answer pairs, including annotated entities and timestamps, with 350k for training and 30k for validation and testing. The dataset can be categorized into simple reasoning (Simple Entity and Simple Time) and complex reasoning (Before/After, First/Last, and Time Join) based on temporal constraints. The TimeQuestions dataset Sharma et al. (2023) has 13.5k manually edited questions and is divided into three parts: training, validation, and testing, containing 7k, 3.2k, and 3.2k questions, respectively. The questions are categorized into four types: Explicit, Implicit, Temporal and Ordinal. Following previous studies, we use two popular evaluation metrics, Hits@1 and Hits@10. More information about datasets and metrics can be found in Appendix A.
We compare four types of baselines: 1) KG embedding-based models including EaE Feng et al. (2020b) and EmbedKGQA Saxena et al. (2020); 2) TKG embedding-based models including CronKGQA Saxena et al. (2021), EntityQR Mavromatis et al. (2022), TMA Liu et al. (2023a), TSQA Shang et al. (2022) and TempoQR Mavromatis et al. (2022); and 3) Language models, including BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and ChatGPT. For another TimeQuestions dataset, we use CronKGQA, TempoQR and TwiRGCN Sharma et al. (2023) for comparison. The implementation details of the baselines are described in Appendix B.
Model Overall Explicit Implicit Temporal Ordinal
CronKGQA 0.462 0.466 0.445 0.511 0.369
TempoQR 0.416 0.465 0.360 0.400 0.349
TwiRGCN(average) 0.605 0.602 0.586 0.641 0.518
TwiRGCN(interval) 0.603 0.599 0.603 0.646 0.494
GenTKGQA 0.584 0.596 0.611 0.563 0.578
Table 2: Hits@1 for different models on TimeQuestions.
@@ 5.2 Main Results
Table 1 reports the performance of all methods on the CronQuestions dataset for various question types. We can observe that GenTKGQA consistently outperforms the baselines in terms of "Overall" performance, and achieves significant improvements of 11.3% in the "Complex" question type and 9.6% in the "Time" answer type on the Hits@1 metric over the second best method. Especially, our model achieves nearly 100% for the "Simple" question type. The possible reason is that simple questions usually involve single facts, GenTKGQA can easily retrieve the relevant facts containing the answer and infer the correct answer through instruction tuning technique.
Furthermore, compared to KG embedding methods, temporal KG embedding methods show significant results on various metrics, thanks to the fact that the temporal information of the TKG is taken into account in the question representation. This is also why KG embedding methods are particularly ineffective for the "Time" answer type. However, most TKG embedding methods treat the QA task as a link prediction, which works for the "Simple" question type containing a single fact compared to the "Complex" question type.
We find that PLMs (BERT, RoBERTa) and LLMs (ChatGPT) have the lowest performance on the TKGQA task without TKG information. This suggests that language models (LM), whether encoded or generated, with a large or small number of parameters, have difficulty answering temporal questions without any relevant context. w/ tkg indicates that LMs use entity/time embeddings or relevant facts from the TKG. Obviously, LMs w/ tkg have significantly better performance, which suggests that the language models have some degree of temporal reasoning capability when relevant TKG information is provided, validating the importance of the subgraph retrieval phase. It is worth noting that ChatGPT w/ tkg still performs weakly in reasoning about complex problem types when providing the facts retrieved in the first stage by GenTKGQA, while our model achieves the best results. This demonstrates the effectiveness of interacting GNN and LM representations in dealing with complex temporal problems. The above findings demonstrate the adequacy of our two motivations for solving complex temporal problems with LLMs. A possible reason for the poor improvement of our model’s Hits@10 metric for the "Entity" type is that the LMs provide irrelevant responses when asked to generate multiple answers.
As shown in Table 2, GenTKGQA still achieves surprising results on the TimeQuestions dataset, especially improving on the "Implicit" and "Ordinal" question types by 1.3% and 11.6%, respectively. Both question types have implied time constraints, similar to the "Complex" question type on the CronQuestions dataset. Overall, compared to the baseline methods, GenTKGQA achieves significant results on the complex question type with different datasets, validating the motivation of our work and the effectiveness of the proposed model. Last but not least, GenTKGQA, as a generative QA model, achieves better results than most traditional extractive QA methods.
@@
Before/ Who held the position of Sociétaire of the
After Comédie-Française before Catherine Hiegel?
[François Jules, position held, Comédie-Française, 1850, 1894]
[Jean Martinelli, position held, Comédie-Française, 1930, 1950] [Lise Delamare] [Yvonne Gaudeau]
[Yvonne Gaudeau, position held, Comédie-Française, 1950, 1986]
[Lise Delamare, position held, Comédie-Française, 1951, 1967]
@@
Time Who was Iowa Governor in Greater German Reich
Join during the World War II?
[Bourke B. H., position held, Governor of Iowa, 1943, 1945] [George A. W.] [George A. W.,
[George A. W., position held, Governor of Iowa, 1939, 1943] Bourke B. H.]
[Robert D. B., position held, Governor of Iowa, 1945, 1949]
@@
5.3 Ablation Study
w/o SR means that we directly perform problem inference without using relevant subgraph information in the model training and inference phases, while w/o SR inference means that we do not provide subgraphs only at the inference. We can observe a sharp decrease in model effectiveness due to the lack of use of subgraph information, which is consistent with the other LMs (Section 5.2). This result shows that current LMs are weak in dealing directly with temporal reasoning problems, validating the importance of the subgraph retrieval module. The w/o SR inference result indicates that GenTKGQA remembers part of the structured knowledge during the training phase and improves the temporal inference performance without providing subgraph information. w/ SR random denotes the random selection of relevant facts involving entities in the question, and the drop in results proves the validity of our first-stage approach.
w/o T-GNN indicates that we directly use the temporal embeddings Lacroix et al. (2020) to represent entities and relations of subgraphs. The slight decrease in the results indicates that the T-GNN is able to perceive the structural information of the TKG. w/o VKI means that we try to remove the virtual knowledge indicators from the input prompt. Model performance degradation shows that indicators can bridge the gap between distinct representations.
Figure 3: Parameter sensitivity on our GenTKGQA.
@@ 5.4 Sensitivity Analysis
We explore the impact of different training data sizes to reason about complex temporal questions. As shown in Figure 3(a), by comparing the Hits@1 metric of several methods for the complex question type, our method consistently outperforms others as the training data expands. In particular, at 20% of the training data, our model outperforms the second best model by 32%, demonstrating that our model has strong inference ability in the case of few-shot samples due to its intrinsic knowledge.
@@
We report the performance changes on the CronQuestions dataset by varying the number of retrieved facts $n$ in Figure 3(b). It can be seen that the model performs poorly with a small number of relevant facts ($n$=3), and there is a slight increase in performance at $n$=10. Fewer facts do not provide sufficient context knowledge, while more facts may introduce noise. Taking this into consideration, we set the hyper-parameter $n$ to 10.
@@ 5.5 Qualitative Results
When providing relevant facts retrieved by GenTKGQA as background knowledge, ChatGPT performs competitively in the simple question type, correctly answering questions with entity or time as the answer. However, it has difficulty answering complex types of questions. For example, in the "Before/After" and "First/Last" question types, ChatGPT struggles to understand the temporal order of the relevant facts and gives incorrect answers. Besides, in the case of the "Time Join" question type, some of the correct answers are missing from the generated responses because ChatGPT does not fully understand the implicit time in the question. On the contrary, GenTKGQA performs well in both simple and complex question types due to the fact that we use a deep manner to incorporate subgraph information into the LLM. However, similar to other LLMs, GenTKGQA randomly generates some irrelevant answers when generating multiple answers, e.g., in the "Simple Entity" question type.
@@ 6 Conclusion
We propose a novel generative framework, GenTKGQA, which guides the LLM in a two-stage manner to handle temporal question answering on TKG. Specifically, at the subgraph retrieval phase, we exploit the LLM’s intrinsic knowledge to mine the temporal constraints and structural links in the temporal questions, which reduces the search space of the subgraphs in both temporal and structural dimensions. We employ the in-context learning approach to complete subgraph retrieval for the entire dataset with a small number of samples. In order to improve the inference performance of the LLM on complex question types, at the answer generation phase, we present the instruction tuning technique to make the open-source LLM truly understand the temporal order and structural dependencies among retrieved facts. Most significantly, we design novel virtual knowledge indicators to establish a bridge between subgraph neural information and text representations. Experimental results show that our framework can effectively utilize the LLM to solve the complex question type of TKGQA task and validate the adequacy of our motivation.
@@
Although the complex temporal question on the CronQuestions dataset involves multiple facts, the inter-entity connection in each fact is single-hop, so the hyper-parameter $k$ of our model is set to 1 to achieve the best results. In fact, the vast majority of current TKGQA datasets involve facts that are single-hop. So, in the future, we will explore more datasets to solve inference for multi-hop complex temporal problems over TKG. In addition, we use the in-context learning approach to prompt the ChatGPT baseline to answer the questions, saving the labor cost of checking whether the answers are correct. However, the design of different templates may result in incomplete consistency with the manual results, but this does not affect the conclusions of this paper. Because the results provided by other works similarly show the poor performance of ChatGPT’s temporal question answering Tan et al. (2023); Chen et al. (2023a).
Ethics Statement
This work presents a novel two-stage framework for the temporal knowledge graph question answering task using large language models. Our experiments use the publicly available datasets and language models from open sources. The dataset is developed to be used for the TKG-based temporal QA task. The language models are used to generate answers to temporal questions with entities or timestamps, which does not involve toxic content. This paper uses the above dataset and models with their initial intention. We believe that this work is consistent with ACL’s ethics policy and presents no potential risk.
Acknowledgements
This work is sponsored in part by the National Natural Science Foundation of China under Grant No. 62025208, and the Xiangjiang Laboratory Fund under Grant No. 22XJ01012.
@@ Appendix A Dataset Statistics and Metrics
We use the CronQuestions and TimeQuestions dataset in our experiments. TimeQuestions was first proposed by Jia et al. (2021), but it provides only static knowledge graphs with temporal attributes, not strictly temporal knowledge graphs. Later, Sharma et al. (2023) expanded this dataset by preprocessing all the contained facts into the temporal knowledge graph format of (subject entity, relationship, object entity, [start time, end time]) and restricting all times to years. Dataset statistics are described in Table 5 and 3, respectively.
@@ Appendix B Baselines and Implementation Details
We use the OpenAI-API [^4] https://platform.openai.com/docs/api-reference (gpt-3.5-turbo-0613555https://platform.openai.com/docs/models/gpt-3-5-turbo) for all ChatGPT-related experiments, including subsequent ChatGPT baselines.
In the subgraph retrieval phase, we use ChatGPT to mine temporal constraints and structural links between entities and add 5 samples to the in-context learning prompt templates, which are presented as Table 7 and 8. We set $k$=1 for the top-$k$ relations. In the answer generation phase, following Lacroix et al. (2020), we select the dimension of entity/relation/time embeddings to 512. For T-GNN, the layer $l$ is set to 1, the linear transformations $\mathbf{W}_{q}$, $\mathbf{W}_{k}$ and $\mathbf{W}_{m}$ are 512×512, and the $\mathbf{m}$ and $\mathbf{b}$ of the MLP layer are $|\mathcal{E}|$. We use the open-source Llama 2-7B Touvron et al. (2023) for instruction tuning and select up to $n$=10 relevant facts as additional knowledge. The linear projection layer $\mathbf{W}_{p}$ is 512×4096. We fine-tune Llama 2-7B using LoRA Hu et al. (2022) with rank 64. The number of epochs is set to 4 and the learning rate is 3e-4. We use the AdamW optimizer Loshchilov and Hutter (2019) with a fixed batch size of 8. We conduct all the experiments with NVIDIA A100 GPUs, and the results of each experiment are averaged over three runs. We will release the source code upon acceptance.
@@
TSQA Shang et al. (2022): TSQA presents a contrastive learning module that improves sensitivity to time relation words.
TwiRGCN Sharma et al. (2023): TwiRGCN is a method for processing TKGQA tasks with the relational graph convolutional network (RGCN).
@@
ChatGPT: To ensure that the output format meets the expected requirements, we use the in-context learning approach to motivate ChatGPT to answer the questions and provide 5 examples in the prompt template. The specific templates are presented in Table 10 of Appendix C. w/o tkg and w/ tkg differ in whether or not question-relevant facts are provided in the input prompts, which are retrieved in the first stage by our proposed GenTKGQA framework.